### v1

# Minimal Models for RNA Simulations

Abstract
The increasing importance of RNA as a prime player in biology can hardly be overstated. It is suspected that the functions of RNA are linked to their structures and dynamics. Many of the problems in RNA, such as folding and RNA-RNA interactions that drive phase separation even in the absence of proteins, require cations. Because experiments alone cannot directly reveal the dynamics of cation-RNA interactions, well calibrated theory and computations are needed to predict how ions control the behavior of RNA. In this review, we outline the development of coarse-grained models at different resolutions with application to phase separation in low complexity sequences. We also describe folding of ribozymes and riboswitches with a focus on the impact of monovalent and divalent cations. We outline major challenges that need to be overcome to simulate complex problems such as assembly of ribosomes.
Department of Chemistry, The University of Texas at Austin, Austin, TX 78712, USA
School of Pharmacy, University of Nottingham, Nottingham, NG7 2RD, UK
Department of Chemistry, University at Buffalo, NY 14260, USA
1 Introduction
Although RNA sometimes referred to as the “dark matter", the spotlight on RNA biology has never been brighter with the award of Nobel prizes in medicine in two consecutive years, one in 2023 for the COVID-19 vaccine based on nucleoside base modifications of RNA 1, and another in 2024 for the role of microRNA in regulating gene expression 2; 3. These are the latest high profile examples in RNA biology which took off after the demonstration that RNA could function as enzymes over forty years ago 4; 5; 6. In the intervening years the functional repertoire of RNA has continuously expanded producing one surprise after another and with no end in sight. As is always the case in science, the discovery of new phenomena usually create a treasure trove of problems that need quantitative explanations, which in this field require both biophysical experiments 7; 8; 9 and computational approaches 10; 11; 12; 13; 14.
The focus of this review is to highlight the utility of coarse-grained (CG) models in the folding of RNA and phase separation in low complexity sequences. Although not a substitute for atomically detailed simulations in water with explicit ions, there is a great need for CG models in understanding the folding of ribozymes (RNA enzymes) as well in the simulations of multiple chains that span a range of time and length scales. Over the last two decades several studies have shown that CG models, beginning with the initial reports on the mechanical force-induced unfolding of a RNA hairpin 10; 15, have been efficacious in making predictions that compare favorably with experiments. The CG models, at different resolutions, have been particularly important in elucidating the role of divalent cations in controlling RNA folding 16; 17; 18.
For the purposes of this article, let us classify CG models into a few categories depending on the resolution of the RNA and the treatment of ions (Figure 1). (i) In the Single Site Interaction (SIS) model, each nucleotide is represented by a bead 19; 20. Electrostatic interactions associated with phosphate groups and the cations are modeled implicitly using the Debye-Hückel (DH) interactions 19; 20; 21. (ii) In the Three Interaction Site (TIS) model 10, each nucleotide has three sites representing phosphate, sugar, and base. Interactions involving charged groups (phosphate and metal ions) are treated explicitly 16 or implicitly 22, depending on the applications. (iii) If the monovalent (such as Na+ or K+) cation concentration exceeds the concentration of divalent cations greatly, as is the case in several folding experiments, the electrostatic interactions associated with monovalent cations could be treated using the DH potential where as divalent cations (Mg2+ or Ca2+) have to be explicitly treated for accurate description of RNA folding. The mixed description of cations when combined with the TIS model for RNA is referred to as the TIS-CAT model 23. (iv) There are models that use higher resolution (six or seven beads per nucleotide) in which charged interactions are either neglected or treated in an ad hoc manner. Most of these models were not designed to investigate RNA folding 24 or RNA-RNA interactions. Due to space limitations, we only discuss models that have made predictions for experiments or compared simulation results directly with experimental findings. Models that have been developed for the purpose of structure predictions25, while important, are beyond the scope of this review.
Figure 1: Coarse-grained models for RNA. (a) Hierarchy of models at different resolution. In the TIS model, each nucleotide consists of three sites representing phosphate (red), sugar (yellow), and base (blue), whereas in the SIS model, a single bead represents an entire nucleotide. (b) Treatment of ions in the TIS model framework. Left: When the solvent contains only monovalent ions, implicit ion treatment based on Debye-Hückel theory is appropriate. Middle: RNA in a solution with monovalent salts and divalent cations such as Mg2+ and Ca2+. Right: In the TIS-CAT model, divalent ions are treated explicitly while monovalent ions are implicit.
2 Single Interaction Site (SIS) Model
The Single Interaction Site (SIS) model represents each nucleotide by single interaction center 19. The SIS model for folded RNA retains chain connectivity and favorable attractive interactions between sites that stabilize the native fold. The initial formulation of the SIS model was primarily used to provide a theoretical framework to predict the outcomes of single molecule pulling experiments in which mechanical forces were to probe the folding landscapes of RNA constructs and ribozymes  26; 27; 28. We first describe applications to force-driven conformational changes in a riboswitch to demonstrate the versatility of the model and set the stage for phase separation studies.
Response of Riboswitches to Mechanical Force: Riboswitches, consiting of an aptamer domain (AD) and an expression platform, contain RNA elements in the untranslated region of mRNA that allosterically regulate translation and transcription in bacteria by binding to metabolites with exquisite specificity 29; 30. Often transcription and riboswitch folding are coupled 31; 32; 33. The functional selectivity is best illustrated using the purine riboswitches as examples. For example, adenine (add) riboswitch 34 activates translation upon binding the metabolite purine, whereas the structurally similar pbuE adenine riboswitch 28 activates transcription in the presence of purine. Both of these are ON riboswitches, which means that translation or transcription is activated only upon binding of the metabolite. In contrast, OFF riboswitches shut down gene expression when the metabolite binds to the aptamer domain. The SAM-III riboswitch binds S-adenosylmethionine (SAM) and represses translation upon binding SAM which sequesters the Shine-Dalgarno ribosome binding sequence. Single molecule force spectroscopy using optical tweezers, performed either by applying constant force to the ends of the RNA or by increasing the force at a constant rate, have been used to assess the characteristics of the intermediates in riboswitches 35; 28. To complement these studies and to describe the structures of the intermediates that the folding landscape of riboswitches could be quantitatively mapped using SIS model 36; 37.
In conventional single-molecular force spectroscopy, performed in the force ramp or force clamp mode, only states that are populated in the accessible force range can be characterized. Experiments performing force pulses during stretch-quench cycles could uncover the hidden states through which RNA folds. Let us start from a fully unfolded riboswitch, prepared at a high stretching force, $f_{s}$, and then quench the force to $f_{q}$ that is less than the critical force needed to unfold for a time period $t_{q}$ to initiate refolding. If $t_{q}$ is less than the folding time RNA would only access partially structured state. By repeating the process of stretching and quenching the force for varying time period ($t_{q}$) folding can be interrupted, and the states that are difficult to capture in a single or step wise quenching of force could be resolved. The SIS model simulations with the metabolite SAM bound showed that a spectrum of states was populated when $t_{q}$ was varied. Importantly, the simulations showed 36 that the refolding time as a function of force is accurately predicted using calculated free energy profile and Kramers rate theory 38; 39.
Phase separation in repeat RNA sequences: More recently, the SIS model has been extended to probe phase separation in low complexity RNA repeat sequences, which cause several neurological and neuromuscular disorders, such as the Huntington disease 40. In vitro experiments 41 showed that the trinucleotide repeat sequences such as (CAG)n and (CUG)n undergo phase separation in which a dense phase coexists with a dilute phase. Importantly, phase separation occurs only if the number of repeats $n$ exceeds a critical value. To determine the mechanism of droplet formation using the SIS model, we adjusted the only parameter in the SIS model, corresponding to the energy scale for base-pair interactions, to reproduce the known structures of a short (CAG)2 duplex. Multichain simulations recapitulate the length and concentration dependence of the phase separation, in quantitative agreement with the in vitro experiments (Figure 2 a and b). The driving force for coacervate formation is the intermolecular base-pair interactions that has both an enthalpic and entropic contribution. Unexpectedly, it was found that once RNA molecules are recruited in the droplets, they undergo large conformational change, from a hairpin-like conformation in isolation to a stretched state, to form an extensive network of intermolecular interactions (Figure 2c). This soft network, in turn, constrains the RNA conformational fluctuation and mobility, which was rationalized using reptation dynamics 42. The SIS model has been adopted to simulate a number of repeat RNA sequences using the LAMMPS, which apparently enhances the computational efficiency 43. Like the previous work, simulations of a number of low complexity RNA sequences showed that the enhanced number of ways of forming Watson-Crick pairs render stability to the dense droplets.
Based on the SIS model, a theory was advanced for anticipating the propensity of RNA to self-associate. It was discovered that the free energy gap separating the ground state and low lying excited states in repeat sequences are small 21. As a result, such sequences can readily unfold, thus exposing a stretch of nucleotides that can engage in inter molecular base pairing 44; 45, resulting in phase separation. In contrast, the free energy gap in heterogeneous sequences in which the 5 and 3 ends are in proximity are sufficiently great that the excited state has negligible population, which prevents unwarranted RNA-RNA interactions.
The single bead representation of RNA has recently been used to investigate phase separations in mixture of RNA and proteins.20; 46; 47 In these models, interactions of RNA residues are parameterized to reproduce condensate formation with proteins, but the capability of such models to fold RNA into complex secondary and tertiary structures has not been demonstrated. Understanding the impact of RNA in influencing phase separation in proteins will also require a consistent model of RNA as well as interactions with ions.
Figure 2: RNA phase separation at the molecular level using coarse-grained simulations. (a, b) The single interaction site model reproduces droplet formation (a) that occurs only when the RNA with $\sim$30 CAG repeats and RNA concentration exceeds critical concentration (b) 48. (c) Snapshots showing incorporation event of a (CAG)47 molecule (red) into a droplet. The dramatic conformational change that occurs can be seen in the accompanying secondary structure diagrams. Left: Initially, the isolated (CAG)47 chain formed a hairpin shape due to self-complementary base pairs. Middle: The RNA is unwound from one end to form new base pairs with other (CAG)47 chains present on the surface of droplet. Right: Once incorporated into the droplet, each RNA chain tends to elongate as formation of inter-chain base pairs dominate due to entropic gain.
3 Three Interaction Site (TIS) Model
In the Three Interaction Site (TIS) model each nucleotide consists of three distinct interaction sites – phosphate (P), sugar (S), and nucleobase (B) (Fig. 1). This allows the TIS model to capture the chemically distinct behavior of these moieties, which is essential for describing the sequence dependent structural dynamics. Since its first application in 2005 10, the TIS model has been widely used to study RNA folding in a variety of contexts. Examples include RNA folding induced by temperature jump 15 or mechanical forces 49, folding pathways shifting with changes in ion concentration on the folding pathways 50, and effects of molecular crowding51, and solvent viscosity 52. The three-bead representation has also been widely adopted in other RNA models 25; 53 and to investigate various aspects of DNA biophysics 54; 55.
Over the years dramatic improvements have been within the TIS model by incorporating the nearest neighbor thermodynamics into the TIS representation 22. This was validated by quantitative comparison with experiments on RNA thermodynamics, such as melting of RNA hairpins and pseudoknots (PKs) (Figure 3a), which are essential building blocks of larger RNA molecules 22; 49. Recent work has focused on developing methods to accurately describe electrostatic interactions associated with monovalent and divalent cations and phosphate groups (Figure 1b). In most CG models of RNA, electrostatic interactions are modeled using the DH potential, which is sufficient for monovalent ions. In fact, for example, the TIS model with DH interactions accurately captured that folding pathways of a PK as a function of monovalent ion concentration, in accord with experiments 50. However, it was a long-standing limitation of CG models that the effects of divalent cations cannot be described accurately using the DH theory. This limitation was overcome in an important study that proposed a way to include both monovalent Na+, K+, Cl-) and divalent ions (Mg2+, Ca2+) in the TIS model explicitly 16. Simulations using this model captures the complicated ion-RNA electrostatic interactions, enabling visualization of how binding of individual cations facilitate self-assembly of RNA molecules.
In a series of papers we illustrated the success of the TIS model with explicit ions by quantitatively reproducing experiments on the folding of the Azoarcus group I intron ribozyme 16; 56; 17. In particular, the simulations reproduced the Mg2+ concentration dependence of folding 16. Furthermore, specific Mg2+ binding sites agreed with the experimentally determined positions of Mg2+ 16; 56 (Figure 3b). This allowed the authors to confidently discuss how individual Mg2+ ions navigate the formation of secondary and tertiary structural motifs in order, which cannot be traced experimentally. The model also captured a subtle difference between Mg2+ and Ca2+. Ca2+ can drive RNA folding to some extent similar to Mg2+, but not enough to reach its enzymatically active conformation due to the slightly larger size of the ionic radius.
It has been shown experimentally that RNA of a certain length with a complex fold often gets caught in kinetic traps and takes a long time to fold or misfold 57. The TIS simulation of the Azoarcus ribozyme has recently been extended to a kinetics study that showed how such multi-step folding and misfolding of the RNA occurs upon sudden changes in solution conditions, such as Mg2+ concentration jump 17. By simulating a series of Mg2+-induced folding events, it has been shown that the fate of the RNA molecules – whether they rapidly fold to the catalytically active state, get trapped in intermediate states, or even misfold – can be determined at a very early stage when the collapsing chain arrangement is driven by Mg2+ ion binding to specific sites (Figure 3c). Notably, the major misfolded state that appeared in the simulation was similar to misfolded structures of Tetrahymena ribozyme recently observed by cryo-electron microscopy 58, giving further confidence to the simulation model.
In an important application of the TIS model, ions (Mg2+, K+, and F-) were included explicitly to investigate the Mg2+-dependent transition of a fluoride riboswitch 59 folding in the presence of K+ and F- ions 60. Kumar and Reddy investigated the folding of the aptamer domain (AD), which led them to propose a novel mechanism for the function of the fluoride riboswitch and F- encapsulation by the cationic pocket composed of three Mg2+ ions. Folding of the AD induces sequential binding of the two Mg2+ ions to the phosphate groups driven in part by a transition from an outer to an inner-shell coordination through dehydration (Fig 4a). The third Mg2+ and F- ion bind to the riboswitch in two steps. Binding of Mg2+ results in trigonal cationic pocket. Subsequently, Mg2+ and F- form a water-mediated ion pair (Mg-water-F).
Let us highlight couple of additional applications of the TIS model. Simulations of the central domain of the bacterial ribosome, performed over a broad range of Mg2+ concentration, reproduced a number of experimental findings quantitatively (Figure 3d) 61. Strikingly, the TIS model simulations predicted not only the multi-pathway folding of G-quadruplex folding but also characterized the structures of the intermediates, thus complementing the experimental results 62. The handful of applications show that the combination of TIS model simulations and experiments could be used to fully determine the folding of large RNAs.
Figure 3: Ion-driven thermodynamics and kinetics of RNA folding driven by ions. (a) Melting profile of BWYV pseudoknot measured in experiment (black)63 at 500 mM NaCl and TIS simulation (red) with implicit monovalent salt at the same concentration. (b) Nucleotide position dependent ion fingerprints in the folding of Azoarcus ribozyme, represented by local concentrations of Mg2+ (blue) and K+ (green). Comparisons with the Mg2+ binding sites resolved in X-ray crystal structure are shown. (c) Left: Kinetic partitioning mechanism in the folding of ribozyme. Differences in partial formations of secondary and tertiary structures at an early stage, driven by specific binding of Mg2+, lead to dramatic difference in the fate – folding to the native structure or misfolding. A major source of misfolding is caused by topological frustration where the position of J8/7 strand (red) relative to neighboring helices are opposite to that in in the native state. Right: The same mechanism of misfolding was recently observed in cryo-EM experiments of a similar ribozyme from Tetrahymena 58. The images were adopted from 17 under the CC-BY license. (d) Left: Conformational transitions driven by Mg2+ in the central domain of bacterial ribosome. Right: Comparison of the high density spots of Mg2+ around the central junction obtained in the simulation (grey) with the experimentally observed Mg2+ positions (red).
4 TIS-CAT
A technical challenge in explicitly treating both divalent and monovalent ions in simulations is that it dramatically decreases the computational efficiency. To overcome this problem the TIS-CAT hybrid model was developed in which monovalent ions are implicitly treated using the DH potential but divalent cations are explicitly included (Figure 1b) 23. The TIS-CAT model is valid when the concentration of monovalent ions (typically $>$50 mM) is higher than the divalent ion concentration, which is normally in the millimolar range in the folding experiments. We developed a comprehensive theoretical model combining the Reference Interaction Site Model (RISM) 64; 65; 66, grounded in liquid state theory, with coarse-grained molecular simulations. The RISM theory enables detailed calculations of the potential of mean force (PMF), which is distance and the ionic concentration-dependent effective potential between divalent ions and the phosphate group (Figure 4a). The isotropic PMF has two distinct minima: one for inner-sphere coordination (direct binding to the phosphate) and another for outer-sphere coordination (binding through water mediation). This dual-mode binding is crucial for accurate prediction of the thermodynamics of ion-induced RNA folding 23.
Simulations using the hybrid model demonstrated that divalent ions, particularly Mg2+, have a profound effect on RNA folding through both inner- and outer-sphere coordination with phosphate groups. In smaller RNAs, such as the Beet Western Yellow Virus (BWYV) pseudoknot, Mg2+ predominantly binds via outer-sphere interactions, where the ion remains hydrated and interacts indirectly with RNA through water molecules. In contrast, in larger RNAs like the 58-nucleotide ribosomal RNA fragment (58-nt rRNA, Figure 4b), there is a significant change toward inner-sphere binding, where Mg2+ directly interacts with phosphate groups after partially losing its hydration shell. This change is due to the increased electrostatic potential in the more densely packed phosphate regions of larger RNA molecules, which necessitates stronger binding to stabilize the folded structure.
The simulations also highlighted the role of bridging interactions on RNA stability 67. These occur when a single divalent ion binds simultaneously to two or more phosphate groups, acting as a bridge that stabilizes the RNA’s tertiary structure. Bridging interactions are particularly important in larger RNAs where multiple phosphate groups are in proximity. The simulations showed that reducing the concentration of divalent ions reduces the likelihood of the bridging interactions, leading to RNA destabilization. For example, in the 58-nt rRNA, specific nucleotides were identified as key binding sites for Mg2+ , with significant accumulation of ions around them 23. These nucleotides are located at structurally important positions, and the Mg2+ ions in these regions are crucial for maintaining the integrity of the folded RNA. This finding is particularly important for understanding the stability of larger RNA molecules, such as ribozymes and ribosomal RNAs, where multiple regions of the RNA come into proximity during folding.
The charge density of divalent cations plays a crucial role in determining their coordination mode with RNA. The high charge density of Mg2+ favors outer-sphere binding in most small and intermediate-sized RNAs because it maintains a stable hydration shell. However, in regions of highly negative electrostatic potential, such as in larger RNA molecules or deep pockets of compact structures, Mg2+ can partially dehydrate and engage in inner-sphere coordination, as in the fluoride riboswitch. In contrast, Ca2+ , which has a lower charge density and thus weaker interaction with its hydration layer, dehydrates more readily and binds directly to phosphate groups in the inner-sphere coordination, even in smaller RNAs. This difference in binding behavior is important for understanding how different divalent ions influence RNA folding and stability. The findings suggest that Mg2+ binds RNA in multiple ways depending on the RNA size and structure, while Ca2+ is more rigid in its preference for direct phosphate binding.
Figure 4: (a) The effective potential between phosphate and Mg2+ used in the TIS-CAT model (red) is a combination of the PMF derived from the RISM theory for the short range (grey dashed line) interaction and the Debye-Hückel theory for the long range (cyan) potential. There are two minima corresponding to the direct (dehydrated) and water-mediated binding of Mg2+ to the phosphate. (b) Preferential interaction coefficient, $\Gamma$, depending on the bulk Mg2+ concentrations as a function of monovalent salt concentration. $\Gamma$ values calculated from TIS-CAT simulations (symbols)23 are in excellent agreement with the experimental data (dashed curves)68. (c) Three distinct conformations of guanidine-II riboswitch from TIS-CAT simulations at different Mg2+ concentrations (upper) and secondary structures indicating the positions of experimental FRET dies (lower) 18. (d) Comparison of smFRET efficiency distribution between the simulations and experiments.18 The images in (c and d) were taken from 18 under the CC-BY license © 2022 Fuks, Falkner, Schwierz and Hengesbach.
Recently, single molecule FRET experiments and simulations using the TIS-CAT model were combined to detect three distinct states that are populated at different Mg2+ concentrations in guanidine-II riboswitch that controls translation (Figure 4c) 18. When guanidine is bound (unbound) translation is turned on (suppressed) 69. Remarkably, the calculated distributions of FRET efficiency as a function Mg2+ using the TIS-CAT model without any adjustable parameter were in excellent agreement with experiments (Figure 4d)18. The TIS-CAT simulations were used to determine the structures of the states that are populated at three values of Mg2+ concentration. Note that unlike in guanidine-II riboswitch it is only upon binding to the metabolite is translation is suppressed in the SAM riboswitch. Nevertheless, the CG models accurately predict the outcomes of experiments, albeit using different resolution for the RNA.
5 Concluding Remarks
Although coarse-grained models for nucleic acids were first introduced only twenty-five years ago 10, a great deal of progress has been made in the intervening years, spurred largely by advances in the experimental fronts (such as single molecule studies, cryo-EM, time-resolved SAXS experiments, and ion-counting experiments). While these developments, some of which are sketched here, are encouraging, frontiers in RNA biology are rapidly moving to problems that invariably require coming to grips with more interactions involving multiple proteins, DNA, and RNA. The prime example is ribosome assembly, which has been extensively studied experimentally 70; 71; 72, thus producing accurate data. Although there are combination of a variety of simulation methods and experiments have provided invaluable insights 73, it is fair to say that computational methods alone are currently not capable of tackling the dynamics of ribosome assembly or similarly complicated problems. We believe that innovations in this context are sorely needed.
Acknowledgments: This work was supported by a grant from the National Science Foundation (CHE 2320256) and the Welch Foundation through the Collie-Welch Chair (F-0019). NH was supported by JST PRESTO (JPMJPR22EA).

### v2

@@
 The increasing importance of RNA as a prime player in biology can hardly be overstated.
-It is suspected that the functions of RNA are linked to their structures and dynamics.
-Many of the problems in RNA, such as folding and RNA-RNA interactions that drive phase separation even in the absence of proteins, require cations.
-Because experiments alone cannot directly reveal the dynamics of cation-RNA interactions, well calibrated theory and computations are needed to predict how ions control the behavior of RNA.
-In this review, we outline the development of coarse-grained models at different resolutions with application to phase separation in low complexity sequences.
-We also describe folding of ribozymes and riboswitches with a focus on the impact of monovalent and divalent cations.
-We outline major challenges that need to be overcome to simulate complex problems such as assembly of ribosomes.
-Department of Chemistry, The University of Texas at Austin, Austin, TX 78712, USA
-School of Pharmacy, University of Nottingham, Nottingham, NG7 2RD, UK
-Department of Chemistry, University at Buffalo, NY 14260, USA
+Problems in RNA, such as folding and RNA-RNA interactions that drive phase separation, require cations.
+Because experiments alone cannot reveal the dynamics of cation-RNA interactions, well calibrated theory and computations are needed to predict how ions control the behavior of RNA.
+The perspective describes the development and use of coarse-grained models at different resolutions.
+We focus on single- and three-interaction site interaction models, in which electrostatic interactions are treated using a combination of explicit and implicit representations.
+Applications to the folding of ribozymes and riboswitches are discussed, with emphasis on the role of monovalent and divalent cations.
+We also discuss phase separation in low complexity sequences.
+Challenges in the simulation of complex problems such as ribosome assembly and RNA chaperones, requiring developments of models for RNA-protein interactions, are pointed out.
+1. Department of Chemistry, The University of Texas at Austin, Austin, TX 78712, USA
+2. School of Pharmacy, University of Nottingham, Nottingham, NG7 2RD, UK
+3. Department of Chemistry, University at Buffalo, NY 14260, USA
 1 Introduction
-Although RNA sometimes referred to as the “dark matter", the spotlight on RNA biology has never been brighter with the award of Nobel prizes in medicine in two consecutive years, one in 2023 for the COVID-19 vaccine based on nucleoside base modifications of RNA 1, and another in 2024 for the role of microRNA in regulating gene expression 2; 3.
-These are the latest high profile examples in RNA biology which took off after the demonstration that RNA could function as enzymes over forty years ago 4; 5; 6.
-In the intervening years the functional repertoire of RNA has continuously expanded producing one surprise after another and with no end in sight.
-As is always the case in science, the discovery of new phenomena usually create a treasure trove of problems that need quantitative explanations, which in this field require both biophysical experiments 7; 8; 9 and computational approaches 10; 11; 12; 13; 14.
-The focus of this review is to highlight the utility of coarse-grained (CG) models in the folding of RNA and phase separation in low complexity sequences.
-Although not a substitute for atomically detailed simulations in water with explicit ions, there is a great need for CG models in understanding the folding of ribozymes (RNA enzymes) as well in the simulations of multiple chains that span a range of time and length scales.
-Over the last two decades several studies have shown that CG models, beginning with the initial reports on the mechanical force-induced unfolding of a RNA hairpin 10; 15, have been efficacious in making predictions that compare favorably with experiments.
-The CG models, at different resolutions, have been particularly important in elucidating the role of divalent cations in controlling RNA folding 16; 17; 18.
+Although RNA sometimes referred to as the “dark matter" of biology, the spotlight on RNA biology has never been brighter.
+The award of Nobel prizes in medicine in two consecutive years, one in 2023 for the vaccines based on nucleoside base modifications of RNA 1, and another in 2024 for the role of microRNA in regulating gene expression 2; 3 highlight the importance of RNA.
+These are the latest high profile examples in RNA biology, which took off after the demonstration that RNA could function as enzymes over forty years ago 4; 5; 6.
+In the intervening years the functional repertoire of RNA has continues to expand producing one surprise after another and with no end in sight.
+As is always the case in science, the discovery of new phenomena create a treasure trove of problems that need quantitative explanations, which in the studies of RNA folding and RNA-RNA interactions require both biophysical experiments 7; 8; 9; 10 and computational approaches 11; 12; 13; 14; 15; 16; 17; 18; 19.
+Here, we focus exclusively on the utility of coarse-grained (CG) models in the folding of RNA and phase separation in low complexity sequences.
+Although not a substitute for atomically detailed simulations in water with explicit ions 20, there is also a need for CG models in understanding the folding of ribozymes (RNA enzymes) as well as in the simulations of multiple chains that span a range of time and length scales.
+hver the last two decades several studies have shown that CG models, beginning with the initial reports on the mechanical force-induced unfolding of a RNA hairpin 11; 21, have been efficacious in making predictions that compare favorably with experiments.
+The CG models, at different resolutions, have been particularly important in elucidating the role of divalent cations in controlling RNA folding 22; 23; 24; 19.
+It is also possible to convert a coarse-grained representation into atomic coordinates for further analysis 25.
 For the purposes of this article, let us classify CG models into a few categories depending on the resolution of the RNA and the treatment of ions (Figure 1).
-(i) In the Single Site Interaction (SIS) model, each nucleotide is represented by a bead 19; 20.
-Electrostatic interactions associated with phosphate groups and the cations are modeled implicitly using the Debye-Hückel (DH) interactions 19; 20; 21.
-(ii) In the Three Interaction Site (TIS) model 10, each nucleotide has three sites representing phosphate, sugar, and base.
-Interactions involving charged groups (phosphate and metal ions) are treated explicitly 16 or implicitly 22, depending on the applications.
-(iii) If the monovalent (such as Na+ or K+) cation concentration exceeds the concentration of divalent cations greatly, as is the case in several folding experiments, the electrostatic interactions associated with monovalent cations could be treated using the DH potential where as divalent cations (Mg2+ or Ca2+) have to be explicitly treated for accurate description of RNA folding.
-The mixed description of cations when combined with the TIS model for RNA is referred to as the TIS-CAT model 23.
-(iv) There are models that use higher resolution (six or seven beads per nucleotide) in which charged interactions are either neglected or treated in an ad hoc manner.
-Most of these models were not designed to investigate RNA folding 24 or RNA-RNA interactions.
-Due to space limitations, we only discuss models that have made predictions for experiments or compared simulation results directly with experimental findings.
-Models that have been developed for the purpose of structure predictions25, while important, are beyond the scope of this review.
+(i) In the Single Interaction Site (SIS) model, each nucleotide is represented by a single bead 26; 27; 28.
+Electrostatic interactions associated with phosphate groups and the cations are modeled implicitly using the Debye-Hückel (DH) interactions 26; 27; 29.
+(ii) In the Three Interaction Site (TIS) model 11, each nucleotide is represented using three sites corresponding to phosphate, sugar, and base.
+Interactions between charged groups (phosphate and metal ions) are treated explicitly 22 or implicitly 30, depending on the applications.
+(iii) Because the concentration of monovalent ions is higher than divalent ions under both physiological and in vitro experimental conditions, we developed a mixed description of the electrostatic interactions and combined it with the TIS model for RNA.
+In the resulting TIS-CAT 31 model, monovalent ions are treated using the DH potential whereas divalent cations (Mg2+ or Ca2+) are explicitly treated for the accurate description of RNA folding.
+(iv) There are models that use higher resolution (six or seven beads per nucleotide) in which stacking and base pairing interactions are treated in an ad hoc manner 18; 32; 33.
+In this review, our main focus is recent development of the first three categories.
+In addition, models that have been developed for the purpose of structure predictions34, while important, are beyond the scope of this review.
+We refer the reader to a more comprehensive review of various RNA simulations and earlier development of CG models 35.
 Figure 1: Coarse-grained models for RNA.
@@
 2 Single Interaction Site (SIS) Model
-The Single Interaction Site (SIS) model represents each nucleotide by single interaction center 19.
+The Single Interaction Site (SIS) model represents each nucleotide by single interaction center 26.
 The SIS model for folded RNA retains chain connectivity and favorable attractive interactions between sites that stabilize the native fold.
-The initial formulation of the SIS model was primarily used to provide a theoretical framework to predict the outcomes of single molecule pulling experiments in which mechanical forces were to probe the folding landscapes of RNA constructs and ribozymes  26; 27; 28.
-We first describe applications to force-driven conformational changes in a riboswitch to demonstrate the versatility of the model and set the stage for phase separation studies.
-Response of Riboswitches to Mechanical Force: Riboswitches, consiting of an aptamer domain (AD) and an expression platform, contain RNA elements in the untranslated region of mRNA that allosterically regulate translation and transcription in bacteria by binding to metabolites with exquisite specificity 29; 30.
-Often transcription and riboswitch folding are coupled 31; 32; 33.
+The initial formulation of the SIS model provided a theoretical framework to predict the outcomes of single molecule pulling experiments in which mechanical forces probed the folding landscapes of RNA constructs and ribozymes  36.
+Response of Riboswitches to Mechanical Force: Riboswitches, consisting of an aptamer domain (AD) and an expression platform, contain RNA elements in the untranslated region of mRNA that allosterically regulate translation and transcription in bacteria by binding to metabolites with exquisite specificity 37; 38.
+Often transcription or translation and riboswitch folding are coupled 39; 40; 41; 42.
 The functional selectivity is best illustrated using the purine riboswitches as examples.
-For example, adenine (add) riboswitch 34 activates translation upon binding the metabolite purine, whereas the structurally similar pbuE adenine riboswitch 28 activates transcription in the presence of purine.
+For example, adenine (add) riboswitch 43 activates translation upon binding the metabolite purine, whereas the structurally similar pbuE adenine riboswitch 36 activates transcription in the presence of purine.
 Both of these are ON riboswitches, which means that translation or transcription is activated only upon binding of the metabolite.
@@
 The SAM-III riboswitch binds S-adenosylmethionine (SAM) and represses translation upon binding SAM which sequesters the Shine-Dalgarno ribosome binding sequence.
-Single molecule force spectroscopy using optical tweezers, performed either by applying constant force to the ends of the RNA or by increasing the force at a constant rate, have been used to assess the characteristics of the intermediates in riboswitches 35; 28.
-To complement these studies and to describe the structures of the intermediates that the folding landscape of riboswitches could be quantitatively mapped using SIS model 36; 37.
-In conventional single-molecular force spectroscopy, performed in the force ramp or force clamp mode, only states that are populated in the accessible force range can be characterized.
-Experiments performing force pulses during stretch-quench cycles could uncover the hidden states through which RNA folds.
-Let us start from a fully unfolded riboswitch, prepared at a high stretching force, $f_{s}$, and then quench the force to $f_{q}$ that is less than the critical force needed to unfold for a time period $t_{q}$ to initiate refolding.
-If $t_{q}$ is less than the folding time RNA would only access partially structured state.
-By repeating the process of stretching and quenching the force for varying time period ($t_{q}$) folding can be interrupted, and the states that are difficult to capture in a single or step wise quenching of force could be resolved.
-The SIS model simulations with the metabolite SAM bound showed that a spectrum of states was populated when $t_{q}$ was varied.
-Importantly, the simulations showed 36 that the refolding time as a function of force is accurately predicted using calculated free energy profile and Kramers rate theory 38; 39.
-Phase separation in repeat RNA sequences: More recently, the SIS model has been extended to probe phase separation in low complexity RNA repeat sequences, which cause several neurological and neuromuscular disorders, such as the Huntington disease 40.
-In vitro experiments 41 showed that the trinucleotide repeat sequences such as (CAG)n and (CUG)n undergo phase separation in which a dense phase coexists with a dilute phase.
-Importantly, phase separation occurs only if the number of repeats $n$ exceeds a critical value.
-To determine the mechanism of droplet formation using the SIS model, we adjusted the only parameter in the SIS model, corresponding to the energy scale for base-pair interactions, to reproduce the known structures of a short (CAG)2 duplex.
+Single molecule force spectroscopy using optical tweezers, performed either by applying constant force to the ends of the RNA or by increasing the force at a constant rate, were used to assess the characteristics of the intermediates in riboswitches 44; 36.
+In conventional single-molecular force spectroscopy, performed in either the force ramp or force clamp mode, only states that are populated in the accessible force range are characterized.
+Experiments performed using force pulses during stretch-quench cycles could uncover the hidden states through which RNA folds.
+Starting from a fully unfolded riboswitch, at a high stretching force, refolding can be initiated by quenching the force to a value less than the critical force needed to unfold for a time period $t_{q}$.
+If $t_{q}$ is less than the folding time, RNA would only access partially structured states.
+By repeating the process of stretching and quenching the force for different values of $t_{q}$ folding can be interrupted, and the states that are difficult to capture in a single or step wise quenching of force could be resolved.
+The SIS model simulations with the metabolite SAM showed that a spectrum of states was populated when $t_{q}$ was varied.
+Importantly, the simulations 45 showed that the refolding time as a function of force is accurately predicted using calculated free energy profile and Kramers rate theory 46.
+Phase separation in repeat RNA sequences: The SIS model has been extended to probe phase separation in low complexity RNA repeat sequences 47.
+In vitro experiments 10 showed that the trinucleotide repeat sequences such as (CAG)n and (CUG)n undergo phase separation in which a dense phase coexists with a dilute phase.
+Interestingly, phase separation occurs only if the number of repeats $n$ exceeds a critical value.
+These findings were explained using the SIS model, with a single unknown energy scale for base-pair interactions, that was adjusted to reproduce the known structures of a short (CAG)2 duplex 16.
 Multichain simulations recapitulate the length and concentration dependence of the phase separation, in quantitative agreement with the in vitro experiments (Figure 2 a and b).
 The driving force for coacervate formation is the intermolecular base-pair interactions that has both an enthalpic and entropic contribution.
-Unexpectedly, it was found that once RNA molecules are recruited in the droplets, they undergo large conformational change, from a hairpin-like conformation in isolation to a stretched state, to form an extensive network of intermolecular interactions (Figure 2c).
-This soft network, in turn, constrains the RNA conformational fluctuation and mobility, which was rationalized using reptation dynamics 42.
-The SIS model has been adopted to simulate a number of repeat RNA sequences using the LAMMPS, which apparently enhances the computational efficiency 43.
+Unexpectedly, it was found that once RNA molecules are recruited in the droplets, they undergo large conformational change, from a hairpin-like conformation in isolation to a nearly stretched state, to form an extensive network of intermolecular interactions (Figure 2c).
+The formation of the soft network, in turn, constrains the RNA conformational fluctuation and mobility, which was rationalized using reptation dynamics 48.
+The SIS model has been adopted to simulate a number of repeat RNA sequences using LAMMPS, which reportedly enhances computational efficiency by speeding up the simulations by a factor of three to five 17.
 Like the previous work, simulations of a number of low complexity RNA sequences showed that the enhanced number of ways of forming Watson-Crick pairs render stability to the dense droplets.
-Based on the SIS model, a theory was advanced for anticipating the propensity of RNA to self-associate.
-It was discovered that the free energy gap separating the ground state and low lying excited states in repeat sequences are small 21.
-As a result, such sequences can readily unfold, thus exposing a stretch of nucleotides that can engage in inter molecular base pairing 44; 45, resulting in phase separation.
-In contrast, the free energy gap in heterogeneous sequences in which the 5 and 3 ends are in proximity are sufficiently great that the excited state has negligible population, which prevents unwarranted RNA-RNA interactions.
-The single bead representation of RNA has recently been used to investigate phase separations in mixture of RNA and proteins.20; 46; 47 In these models, interactions of RNA residues are parameterized to reproduce condensate formation with proteins, but the capability of such models to fold RNA into complex secondary and tertiary structures has not been demonstrated.
-Understanding the impact of RNA in influencing phase separation in proteins will also require a consistent model of RNA as well as interactions with ions.
-Figure 2: RNA phase separation at the molecular level using coarse-grained simulations.
-(a, b) The single interaction site model reproduces droplet formation (a) that occurs only when the RNA with $\sim$30 CAG repeats and RNA concentration exceeds critical concentration (b) 48.
+Based on the SIS model, a theory was advanced for predicting the propensity of RNA to self-associate.
+It was discovered that the free energy gap separating the ground state and low lying excited states in repeat sequences are small relative to $k_{B}T$ 29; 49.
+As a result, such sequences can readily unfold, thus exposing a stretch of nucleotides that can engage in inter molecular base pairing 50; 51, resulting in phase separation.
+In contrast, the free energy gap in heterogeneous sequences in which the 5 and 3 ends are in proximity are sufficiently large 29 that the excited state has negligible population, which prevents unwarranted RNA-RNA interactions.
+The single bead representation of RNA has recently been used to investigate phase separations in mixture of RNA and proteins.27; 52; 53 In these models, interactions of RNA residues are parameterized to reproduce condensate formation with proteins, but the capability of such models to fold RNA into complex secondary and tertiary structures has not been demonstrated.
+Understanding the impact of RNA in influencing phase separation in proteins will also require a consistent models of RNA as well as interactions with ions.
+Figure 2: RNA phase separation using coarse-grained simulations.
+(a, b) The single interaction site model reproduces droplet formation (a) that occurs only when the RNA with $\sim$30 CAG repeats and RNA concentration exceeds the critical concentration needed to cause phase separation (LLPS).
+(b) 16.
 (c) Snapshots showing incorporation event of a (CAG)47 molecule (red) into a droplet.
-The dramatic conformational change that occurs can be seen in the accompanying secondary structure diagrams.
+Dramatic conformational changes occurs as can be seen in the accompanying secondary structure diagrams.
 Left: Initially, the isolated (CAG)47 chain formed a hairpin shape due to self-complementary base pairs.
@@
 3 Three Interaction Site (TIS) Model
-In the Three Interaction Site (TIS) model each nucleotide consists of three distinct interaction sites – phosphate (P), sugar (S), and nucleobase (B) (Fig.
+In the Three Interaction Site (TIS) model, each nucleotide consists of three distinct interaction sites – phosphate (P), sugar (S), and nucleobase (B) (Fig.
 1).
-This allows the TIS model to capture the chemically distinct behavior of these moieties, which is essential for describing the sequence dependent structural dynamics.
-Since its first application in 2005 10, the TIS model has been widely used to study RNA folding in a variety of contexts.
-Examples include RNA folding induced by temperature jump 15 or mechanical forces 49, folding pathways shifting with changes in ion concentration on the folding pathways 50, and effects of molecular crowding51, and solvent viscosity 52.
-The three-bead representation has also been widely adopted in other RNA models 25; 53 and to investigate various aspects of DNA biophysics 54; 55.
-Over the years dramatic improvements have been within the TIS model by incorporating the nearest neighbor thermodynamics into the TIS representation 22.
-This was validated by quantitative comparison with experiments on RNA thermodynamics, such as melting of RNA hairpins and pseudoknots (PKs) (Figure 3a), which are essential building blocks of larger RNA molecules 22; 49.
+This allows the TIS model to capture the chemically distinct behavior of these moieties.
+Since its first application in 2005 11, the TIS model has been widely used to study RNA folding in a variety of contexts.
+Examples include RNA folding induced by temperature jump 21 or mechanical forces 54, folding pathways shifting with changes in ion concentration on the folding pathways 55, and effects of molecular crowding56, and solvent viscosity 57.
+The three-bead representation has also been widely adopted in other RNA models 34; 58 and to investigate various aspects of DNA biophysics 59; 60.
+Over the years dramatic improvements have been within the TIS model by incorporating the nearest neighbor thermodynamics into the TIS representation 30.
+This was validated by quantitative comparison with experiments on RNA thermodynamics, such as melting of RNA hairpins and pseudoknots (PKs) (Figure 3a), which are essential building blocks of larger RNA molecules 30; 54.
 Recent work has focused on developing methods to accurately describe electrostatic interactions associated with monovalent and divalent cations and phosphate groups (Figure 1b).
-In most CG models of RNA, electrostatic interactions are modeled using the DH potential, which is sufficient for monovalent ions.
-In fact, for example, the TIS model with DH interactions accurately captured that folding pathways of a PK as a function of monovalent ion concentration, in accord with experiments 50.
-However, it was a long-standing limitation of CG models that the effects of divalent cations cannot be described accurately using the DH theory.
-This limitation was overcome in an important study that proposed a way to include both monovalent Na+, K+, Cl-) and divalent ions (Mg2+, Ca2+) in the TIS model explicitly 16.
-Simulations using this model captures the complicated ion-RNA electrostatic interactions, enabling visualization of how binding of individual cations facilitate self-assembly of RNA molecules.
-In a series of papers we illustrated the success of the TIS model with explicit ions by quantitatively reproducing experiments on the folding of the Azoarcus group I intron ribozyme 16; 56; 17.
-In particular, the simulations reproduced the Mg2+ concentration dependence of folding 16.
-Furthermore, specific Mg2+ binding sites agreed with the experimentally determined positions of Mg2+ 16; 56 (Figure 3b).
-This allowed the authors to confidently discuss how individual Mg2+ ions navigate the formation of secondary and tertiary structural motifs in order, which cannot be traced experimentally.
-The model also captured a subtle difference between Mg2+ and Ca2+.
-Ca2+ can drive RNA folding to some extent similar to Mg2+, but not enough to reach its enzymatically active conformation due to the slightly larger size of the ionic radius.
-It has been shown experimentally that RNA of a certain length with a complex fold often gets caught in kinetic traps and takes a long time to fold or misfold 57.
-The TIS simulation of the Azoarcus ribozyme has recently been extended to a kinetics study that showed how such multi-step folding and misfolding of the RNA occurs upon sudden changes in solution conditions, such as Mg2+ concentration jump 17.
-By simulating a series of Mg2+-induced folding events, it has been shown that the fate of the RNA molecules – whether they rapidly fold to the catalytically active state, get trapped in intermediate states, or even misfold – can be determined at a very early stage when the collapsing chain arrangement is driven by Mg2+ ion binding to specific sites (Figure 3c).
-Notably, the major misfolded state that appeared in the simulation was similar to misfolded structures of Tetrahymena ribozyme recently observed by cryo-electron microscopy 58, giving further confidence to the simulation model.
-In an important application of the TIS model, ions (Mg2+, K+, and F-) were included explicitly to investigate the Mg2+-dependent transition of a fluoride riboswitch 59 folding in the presence of K+ and F- ions 60.
-Kumar and Reddy investigated the folding of the aptamer domain (AD), which led them to propose a novel mechanism for the function of the fluoride riboswitch and F- encapsulation by the cationic pocket composed of three Mg2+ ions.
-Folding of the AD induces sequential binding of the two Mg2+ ions to the phosphate groups driven in part by a transition from an outer to an inner-shell coordination through dehydration (Fig 4a).
+In most CG models, electrostatic interactions are modeled using the DH potential, which is sufficient for monovalent ions.
+For example, the TIS model with DH interactions accurately captured that folding pathways of a PK as a function of monovalent ion concentration, in accord with experiments 55.
+However, a limitation of such CG models is that the effects of divalent cations could not be described accurately using the DH theory.
+This limitation was overcome in an important study that proposed a way to include both monovalent (Na+, K+, Cl-) and divalent ions (Mg2+, Ca2+) in the TIS model explicitly 22.
+Simulations using this model capture the ion-RNA electrostatic interactions, enabling visualization of how binding of individual cations facilitate the self-assembly of RNA molecules 22; 61.
+In a series of papers, the success of the TIS model with explicit ions has been demonstrated by quantitatively reproducing experiments on the folding of the Azoarcus group I intron ribozyme 22; 62; 23.The simulations reproduced the Mg2+ concentration dependence of folding 22.
+Furthermore, specific Mg2+ binding sites agreed with the experimentally determined positions of Mg2+ 22; 62 (Figure 3b).
+Interestingly, the predictive power of various other methods on Mg2+ ion distributions on the surface of RNA molecules was evaluated at the recent CASP16, reflecting the importance and difficulty of the problem.
+In the TIS simulations, the trajectories were dissected to predict how individual Mg2+ ions facilitate the formation of secondary and tertiary structural motifs, which cannot be tracked experimentally.
+The simulations also showed that subtle difference between Mg2+ and Ca2+.
+Ca2+ can drive RNA folding but not enough to reach the enzymatically active conformation  22.
+Theoretical arguments and experiments have shown that RNA with a complex fold are kinetically trapped in native-like misfolded states, thus requiring considerable time to fold 63.
+The TIS kinetic simulations showed that folds Azoarcus ribozyme folds by a multi-step process involving a direct path to the native state as well as trapping in misfolded metastable native-like states upon a jump in the Mg2+ concentration 23.
+Analyses of a series of Mg2+-induced folding events, revealed that the fate of the RNA molecules – whether they rapidly fold to the catalytically active state, get trapped in intermediate states, or even misfold – occur at the very early stage coinciding with the chain collapse arrangement that is driven by Mg2+ ion binding to specific sites (Figure 3c).
+Notably, the major misfolded state that was identified in the simulation was similar to the misfolded structures of Tetrahymena ribozyme resolved by cryo-electron microscopy 64.
+In an important application of the TIS model, ions (Mg2+, K+, and F-) were included explicitly to investigate the Mg2+-dependent transition of a fluoride riboswitch folding in the presence of K+ and F- ions 65.
+Kumar and Reddy investigated the folding of the aptamer domain (AD), which led them to propose a novel mechanism for the function of the fluoride riboswitch and F- encapsulation by the cationic pocket composed of three Mg2+ ions65.
+Folding of the AD induces sequential binding of two Mg2+ ions to the phosphate groups driven in part by a transition from an outer to an inner-shell coordination through dehydration (Fig 4a).
 The third Mg2+ and F- ion bind to the riboswitch in two steps.
-Binding of Mg2+ results in trigonal cationic pocket.
+Binding of the Mg2+ results in trigonal cationic pocket.
 Subsequently, Mg2+ and F- form a water-mediated ion pair (Mg-water-F).
-Let us highlight couple of additional applications of the TIS model.
-Simulations of the central domain of the bacterial ribosome, performed over a broad range of Mg2+ concentration, reproduced a number of experimental findings quantitatively (Figure 3d) 61.
-Strikingly, the TIS model simulations predicted not only the multi-pathway folding of G-quadruplex folding but also characterized the structures of the intermediates, thus complementing the experimental results 62.
-The handful of applications show that the combination of TIS model simulations and experiments could be used to fully determine the folding of large RNAs.
-Figure 3: Ion-driven thermodynamics and kinetics of RNA folding driven by ions.
-(a) Melting profile of BWYV pseudoknot measured in experiment (black)63 at 500 mM NaCl and TIS simulation (red) with implicit monovalent salt at the same concentration.
+Let us highlight a couple of additional applications of the TIS model.
+Simulations of the central domain of the bacterial ribosome, performed over a broad range of Mg2+ concentrations, reproduced a number of experimental findings quantitatively (Figure 3d) 61.
+Strikingly, the TIS model simulations predicted not only the multi-pathway folding of G-quadruplex folding but also characterized the structures of the intermediates, thus complementing the experimental results 66.
+The handful of applications show that the combination of TIS model simulations and experiments could be used to fully determine the folding of large RNAs, which currently cannot be achieved by other methods.
+Figure 3: Ion-driven thermodynamics and kinetics of RNA folding.
+(a) Experimental melting profile of BWYV pseudoknot (black)67 at 500 mM NaCl and TIS simulation (red) with implicit monovalent salt at the same concentration.
 (b) Nucleotide position dependent ion fingerprints in the folding of Azoarcus ribozyme, represented by local concentrations of Mg2+ (blue) and K+ (green).
@@
 (c) Left: Kinetic partitioning mechanism in the folding of ribozyme.
-Differences in partial formations of secondary and tertiary structures at an early stage, driven by specific binding of Mg2+, lead to dramatic difference in the fate – folding to the native structure or misfolding.
-A major source of misfolding is caused by topological frustration where the position of J8/7 strand (red) relative to neighboring helices are opposite to that in in the native state.
-Right: The same mechanism of misfolding was recently observed in cryo-EM experiments of a similar ribozyme from Tetrahymena 58.
-The images were adopted from 17 under the CC-BY license.
+Differences in partial formations of secondary and tertiary structures at an early stage, driven by specific binding of Mg2+, lead to dramatic difference in the fate – folding to the native structure or native metastable states.
+A major source of misfolding is due to topological frustration where the position of J8/7 strand (red) relative to neighboring helices are opposite to that in in the native state.
+Right: The same mechanism of misfolding was recently observed in cryo-EM experiments of a similar ribozyme from Tetrahymena 64.
+The images were adopted from 23 under the CC-BY license.
 (d) Left: Conformational transitions driven by Mg2+ in the central domain of bacterial ribosome.
@@
 4 TIS-CAT
-A technical challenge in explicitly treating both divalent and monovalent ions in simulations is that it dramatically decreases the computational efficiency.
-To overcome this problem the TIS-CAT hybrid model was developed in which monovalent ions are implicitly treated using the DH potential but divalent cations are explicitly included (Figure 1b) 23.
+A technical challenge in explicitly treating both divalent and monovalent ions in simulations is that it dramatically increases the computational cost.
+For instance, simulating a 35-nucleotide-long pseudoknot RNA requires a system comprising 3,423 beads, 3,303 of which are for K+ and Cl- ions.
+Consequently, the majority of the simulation time is spent computing interactions between monovalent ions.
+To overcome this problem, the TIS-CAT hybrid model was developed in which monovalent ions are implicitly treated using the DH potential but divalent cations are explicitly included (Figure 1b) 31.
 The TIS-CAT model is valid when the concentration of monovalent ions (typically $>$50 mM) is higher than the divalent ion concentration, which is normally in the millimolar range in the folding experiments.
-We developed a comprehensive theoretical model combining the Reference Interaction Site Model (RISM) 64; 65; 66, grounded in liquid state theory, with coarse-grained molecular simulations.
-The RISM theory enables detailed calculations of the potential of mean force (PMF), which is distance and the ionic concentration-dependent effective potential between divalent ions and the phosphate group (Figure 4a).
+In the TIS-CAT model, we used the Reference Interaction Site Model (RISM) 68, grounded in liquid state theory in conjunction with coarse-grained molecular simulations.
+The RISM theory enables the calculations of the potential of mean force (PMF), which is the effective potential between divalent ions and the phosphate group distance that depends on the ion concentration (Figure 4a).
 The isotropic PMF has two distinct minima: one for inner-sphere coordination (direct binding to the phosphate) and another for outer-sphere coordination (binding through water mediation).
-This dual-mode binding is crucial for accurate prediction of the thermodynamics of ion-induced RNA folding 23.
+The dual-mode binding is crucial for accurate prediction of the thermodynamics of ion-induced RNA folding 31.
 Simulations using the hybrid model demonstrated that divalent ions, particularly Mg2+, have a profound effect on RNA folding through both inner- and outer-sphere coordination with phosphate groups.
 In smaller RNAs, such as the Beet Western Yellow Virus (BWYV) pseudoknot, Mg2+ predominantly binds via outer-sphere interactions, where the ion remains hydrated and interacts indirectly with RNA through water molecules.
-In contrast, in larger RNAs like the 58-nucleotide ribosomal RNA fragment (58-nt rRNA, Figure 4b), there is a significant change toward inner-sphere binding, where Mg2+ directly interacts with phosphate groups after partially losing its hydration shell.
+In contrast, in the 58-nucleotide ribosomal RNA fragment (58-nt rRNA, Figure 4b), there is a significant shift toward inner-sphere binding, where Mg2+ directly interacts with phosphate groups after partially losing its hydration shell69.
 This change is due to the increased electrostatic potential in the more densely packed phosphate regions of larger RNA molecules, which necessitates stronger binding to stabilize the folded structure.
-The simulations also highlighted the role of bridging interactions on RNA stability 67.
+The simulations also highlighted the role of bridging interactions on RNA stability 69.
 These occur when a single divalent ion binds simultaneously to two or more phosphate groups, acting as a bridge that stabilizes the RNA’s tertiary structure.
 Bridging interactions are particularly important in larger RNAs where multiple phosphate groups are in proximity.
-The simulations showed that reducing the concentration of divalent ions reduces the likelihood of the bridging interactions, leading to RNA destabilization.
-For example, in the 58-nt rRNA, specific nucleotides were identified as key binding sites for Mg2+ , with significant accumulation of ions around them 23.
+The simulations showed that lowering the concentration of divalent ions reduces the likelihood of bridging interactions, leading to RNA destabilization.
+Individual nucleotides were identified as either forming or not forming interactions with Mg2+, even in early folding or intermediate structures, depending on the bulk Mg2+ concentration.
+For example, in the 58-nt rRNA, specific nucleotides were identified as key binding sites for Mg2+, with significant accumulation of ions around them 31.
 These nucleotides are located at structurally important positions, and the Mg2+ ions in these regions are crucial for maintaining the integrity of the folded RNA.
@@
 The charge density of divalent cations plays a crucial role in determining their coordination mode with RNA.
-The high charge density of Mg2+ favors outer-sphere binding in most small and intermediate-sized RNAs because it maintains a stable hydration shell.
+The high charge density of Mg2+ favors outer-sphere binding in most small and intermediate-sized RNAs because it maintains a stable hydration shell 69.
 However, in regions of highly negative electrostatic potential, such as in larger RNA molecules or deep pockets of compact structures, Mg2+ can partially dehydrate and engage in inner-sphere coordination, as in the fluoride riboswitch.
@@
 (b) Preferential interaction coefficient, $\Gamma$, depending on the bulk Mg2+ concentrations as a function of monovalent salt concentration.
-$\Gamma$ values calculated from TIS-CAT simulations (symbols)23 are in excellent agreement with the experimental data (dashed curves)68.
-(c) Three distinct conformations of guanidine-II riboswitch from TIS-CAT simulations at different Mg2+ concentrations (upper) and secondary structures indicating the positions of experimental FRET dies (lower) 18.
-(d) Comparison of smFRET efficiency distribution between the simulations and experiments.18 The images in (c and d) were taken from 18 under the CC-BY license © 2022 Fuks, Falkner, Schwierz and Hengesbach.
-Recently, single molecule FRET experiments and simulations using the TIS-CAT model were combined to detect three distinct states that are populated at different Mg2+ concentrations in guanidine-II riboswitch that controls translation (Figure 4c) 18.
-When guanidine is bound (unbound) translation is turned on (suppressed) 69.
-Remarkably, the calculated distributions of FRET efficiency as a function Mg2+ using the TIS-CAT model without any adjustable parameter were in excellent agreement with experiments (Figure 4d)18.
+$\Gamma$ values calculated from TIS-CAT simulations (symbols)31 are in excellent agreement with the experimental data (dashed curves)70.
+(c) Three distinct conformations of guanidine-II riboswitch from TIS-CAT simulations at different Mg2+ concentrations (upper) and secondary structures indicating the positions of experimental FRET dies (lower) 24.
+(d) Comparison of smFRET efficiency distribution between the simulations and experiments.24 The images in (c and d) were taken from 24 under the CC-BY license © 2022 Fuks, Falkner, Schwierz and Hengesbach.
+Recently, single molecule FRET experiments and simulations using the TIS-CAT model were combined to detect three distinct states that are populated at different Mg2+ concentrations in guanidine-II riboswitch that controls translation (Figure 4c) 24.
+When guanidine is bound (unbound) translation is turned on (suppressed) 71.
+Remarkably, the calculated distributions of FRET efficiency as a function of Mg2+ concentration using the TIS-CAT model without any adjustable parameter were in excellent agreement with experiments (Figure 4d)24.
 The TIS-CAT simulations were used to determine the structures of the states that are populated at three values of Mg2+ concentration.
-Note that unlike in guanidine-II riboswitch it is only upon binding to the metabolite is translation is suppressed in the SAM riboswitch.
+Note that unlike the guanidine II riboswitch, the SAM riboswitch does not suppress translation until it binds to a metabolite.
 Nevertheless, the CG models accurately predict the outcomes of experiments, albeit using different resolution for the RNA.
+The oxRNA model, which is also based on TIS representation of RNA, has been used to calculate the melting profiles of hairpins and the MMTV pseudoknot 19; 72.
+The initial version 19 of oxRNA did not account for electrostatic interactions.
+As a result, ion dependence of RNA folding was not investigated.
+Nevertheless, the predictions of the two melting temperatures in the MMTV pseudoknot at 1M salt were reasonably accurate.
+Subsequently 72, electrostatic interactions were incorporated using the DH potential.
+Application to supercoiling of double stranded RNA produced good agreement with experiments.
 5 Concluding Remarks
-Although coarse-grained models for nucleic acids were first introduced only twenty-five years ago 10, a great deal of progress has been made in the intervening years, spurred largely by advances in the experimental fronts (such as single molecule studies, cryo-EM, time-resolved SAXS experiments, and ion-counting experiments).
-While these developments, some of which are sketched here, are encouraging, frontiers in RNA biology are rapidly moving to problems that invariably require coming to grips with more interactions involving multiple proteins, DNA, and RNA.
-The prime example is ribosome assembly, which has been extensively studied experimentally 70; 71; 72, thus producing accurate data.
-Although there are combination of a variety of simulation methods and experiments have provided invaluable insights 73, it is fair to say that computational methods alone are currently not capable of tackling the dynamics of ribosome assembly or similarly complicated problems.
-We believe that innovations in this context are sorely needed.
+Since the introduction of coarse-grained models to study RNA folding mechanisms twenty years ago 11, a great deal of progress has been made, spurred largely by advances in the experimental frontiers (such as single molecule studies, cryo-EM, time-resolved SAXS and ion-counting experiments).
+Although these developments, some of which are sketched here, are encouraging, the frontiers of RNA biology are rapidly moving to problems that invariably require coming to grips with complex systems involving multiple proteins, DNA, and RNA.
+To address such challenges, there is a clear need for more sophisticated coarse-grained models capable of accurately describing the formation of multi-component complexes, such as RNA-protein assemblies, while remaining computationally applicable to large-scale systems.
+One example is ribosome assembly, which has been extensively studied experimentally 73; 74; 75.
+A normal mode analysis using elastic network representation provided insights into the dynamics of ribosomes in the ground state 76, and a single-bead coarse-grained simulation captured certain motions beyond the ground state 77.
+Nevertheless, it is fair to say that computational methods alone are currently not capable of tackling the dynamics of ribosome assembly or spliceosomes.
 Acknowledgments: This work was supported by a grant from the National Science Foundation (CHE 2320256) and the Welch Foundation through the Collie-Welch Chair (F-0019).
 NH was supported by JST PRESTO (JPMJPR22EA).
+We thank Balaka Mondal for proofreading the manuscript and the two anonymous referees who provided valuable feedback to help us improve it.